The claim is that the proportion of adults who smoked a cigarette in the past week is less than 0.35​, and the sample statistics include nequals=1562 subjects with 594 saying that they smoked a cigarette in the past week.
Find the value of the test statistic.
Thanks!

Answer by stanbon(75887)

The claim is that the proportion of adults who smoked a cigarette in the past week is less than 0.35​, and the sample statistics include n =1562 subjects with 594 saying that they smoked a cigarette in the past week.
Find the value of the test statistic.
----
Sample proportion = 594/1562 = 0.38
Test Stat:: z(0.38) = (0.38-0.35)/sqrt(0.35*0.65/1562) = 2.4858
-----
Cheers,
Stan H.
--------------